The authors of comic books say that it takes a lot of time to conceive of, illustrate and write a comic book. ( on average, 40 days). Pixton Books wants to determine a minimum sample size needed to test the average time;  given a stdev = 15.23 days and a confidence level of 95%, find the smallest acceptable sample size;  they want to be within 2.9 days.   Be sure to round appropriately.
Answer #1

We have to find sample size n

Given ,

Standard deviation = = 15.23 days

Margin of Error = E = 2.9 days.

Formula :

Where , is critical value at given confidence level.

Confidence level = 95%

Significance level = = 1-0.95 = 0.05 , /2 = 0.05/2 = 0.025

So, = = 1.96 ,     { Using excel function , =NORMSINV(1-0.025) = 1.96 }

So sample size ( n ) is ,

( Round up to next integer )

Minimum required sample size is n = 106